# Value-oriented forecast reconciliation for renewables in electricity markets

### v1

@@ 3.2.2 The Derivation of Cost Allocation
Let $\tilde{y}_{t+k,\texttt{sum}|t}$ be the offer by the PM for the trading hour $t+k$, and ${\hat{y}_{t+k,1|t},\cdots,\hat{y}_{t+k,m|t}}$ be a set of $m$ independent offers. Then, the extra profit is defined as
$R=\sum_{i=1}^{m}c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}}),$
where $c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})$ is the cost for the aggregate offer $\tilde{y}_{t+k,\texttt{sum}|t}$ post-reconciliation, and $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$ is the cost for the independent offer using the base forecast $\hat{y}_{t+k,i|t}$.
@@
$\sum_{i=1}^{m}c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})=(1-w)\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})+wc_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}}).$
@@
$c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})\leq\sum_{i=1}^{m}c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}),$ (24)
where the inequality results from Theorem 3.1. Therefore, it can be seen that the designed cost allocation rule is not efficient, unless $w=1$, as we have
$\sum_{i=1}^{m}c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-\sum_{i=1}^{m}c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq R.$ (25)
In other words, the total extra profit allocated to WPPs is lower than the extra profit generated by aggregated trading. Specifically, the remaining extra profit is assigned to the PM as the payoff for their coordination efforts, denoted as $R^{\texttt{pm}}$, i.e.,
$R^{\texttt{pm}}=(1-w)\left[\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})-c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})\right].$ (26)
@@ 3.3 Model Training via Empirical Risk Minimization
As outlined in problem (21), we compute the parameters of the reconciliation model using the Nash bargaining approach. Nonetheless, neither $\mathbb{E}\left[c_{t+k,i}^{(\texttt{AG})}(f_{\texttt{RE}}(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t};\bm{\theta}),\mathbf{y}_{t+k})\right]$ nor $\mathbb{E}\left[c_{t+k}(\hat{y}_{t+k,i},y_{t+k,i})\right]$ is available. Following Ban and Rudin, (2019), we employ the sample average approximation method to approximate the expectation components in (21) and apply the empirical risk minimization algorithm for parameter estimation. As a result, we convert problem (21) into a constrained learning problem. With a set of historical data $\{\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t},\mathbf{y}_{t+k}\}_{t=1,2,\cdots,T}$, we can rewrite (21) as
$\underset{\bm{\theta}}{\min}$ $\,\quad\sum_{i}^{m}-\log\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]\right],$ (27a)
s.t. $\quad\,\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i},y_{t+k,i})\right]\leq 0,\ \forall i=1,\cdots,m,$ (27b)
Unlike Ban and Rudin, (2019), which employs linear models for data-driven decisions, we leverage neural networks to model the combination function $g$.
Chamon and Ribeiro, (2020) demonstrated that, in theory, learning with constraints is no more difficult than learning without them, and empirical risk minimization is an appropriate algorithm for such learning. Specifically, they proposed a primal-dual gradient algorithm for constrained learning. Here, we adapt their approach and derive the empirical Lagrangian function of problem (27).
$\begin{split}L(\bm{\theta},\bm{\mu})&=-\sum_{i}^{m}\log\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]\right]\\ &+\sum_{i}^{m}\mu_{i}\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+},\end{split}$ (28)
@@
$\nabla_{\mu_{i}}=\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+},$
where $\tilde{\mathbf{y}}_{t+k|t}=\mathbf{S}g(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t};\bm{\theta})$. Hence, the main challenge of problem (29) lies in the inner minimization. However, abundant evidence has shown that the gradient descent algorithm can learn good parameters for neural networks (Shalev-Shwartz and Ben-David,, 2014). Subsequently, we can iteratively minimize the Lagrangian with respect to $\bm{\theta}$ while keeping $\bm{\mu}$ constant, followed by adjusting the dual variables using the obtained minimizer.
@@
     $\mu_{i}^{\eta}\leftarrow\mu_{i}^{\eta-1}+\nu_{i}\left[\frac{1}{|\mathcal{T}_{b}|}\sum_{t\in\mathcal{T}_{b}}\left[c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+}$
@@
We now analyze the convexity of the Lagrangian function. The convexity of the second component of $L(\bm{\theta},\bm{\mu})$ is straightforward, as $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is convex in $\tilde{\mathbf{y}}_{t+k|t}$. For the first component of $L(\bm{\theta},\bm{\mu})$, we note that $-\log(\cdot)$ is convex and decreasing, while $-c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is concave in $\tilde{\mathbf{y}}_{t+k|t}$. Since the combination of a convex, non-increasing function with a concave function preserves convexity (Boyd and Vandenberghe,, 2004), it follows that the first term in $L(\bm{\theta},\bm{\mu})$ is also convex. Therefore, the Lagrangian function $L(\bm{\theta},\bm{\mu})$ is convex in $\tilde{\mathbf{y}}_{t+k|t}$, enabling the use of the stochastic gradient descent algorithm to update $\bm{\theta}$ (Shalev-Shwartz and Ben-David,, 2014). Using a batch of samples $\{\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t},\mathbf{y}_{t+k}\}_{t\in\mathcal{T}_{b}}$ where $\mathcal{T}_{b}$ represents the set of indices, we write the Lagrangian of batch samples as
$\begin{split}L^{(b)}(\bm{\theta},\bm{\mu})&=-\sum_{i}^{m}\log\left[\frac{1}{|\mathcal{T}_{b}|}\sum_{t\in\mathcal{T}_{b}}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]\right]\\ &+\sum_{i}^{m}\mu_{i}\left[\frac{1}{|\mathcal{T}_{b}|}\sum_{t\in\mathcal{T}_{b}}\left[c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+},\end{split}$ (30)
where $|\mathcal{T}_{b}|$ represents the cardinality of $\mathcal{T}_{b}$. Denoting the parameter estimates at the $\eta$ iteration as $\bm{\mu}^{\eta}$ and $\bm{\theta}^{\eta}$, the iterative learning procedure is summarized in Algorithm 1.
@@ 4 Simulations
In this section, we present several illustrative examples to validate the proposed approach using data from the Wind Toolkit [^2] https://www.nrel.gov/grid/wind-toolkit.html. Specifically, we have collected hourly datasets from four nearby wind farms, with capacities of 1.7496 MW, 2.9646 MW, 3.3777 MW, and 2.5272 MW, covering the period from 2007 to 2008. In particular, we follow the framework of Muñoz et al., (2023), which examines an hourly online forward market followed by a balancing market with a dual-price settlement mechanism. The lead time, $k$, for both the base forecast and forecast reconciliation process is set to 1. The dataset is split into two segments: the first 80% is used for model training, while the remaining 20% is reserved for out-of-sample forecasting validation. In the following, we will introduce the setup and present the simulation results.
@@ 4.1 Experimental Setups
In our simulation studies, we investigate a simplified scenario where $\pi_{t+k}^{\texttt{F}}$ is fixed at 25 €/MWh, $\psi^{+}_{t+k}$ is set to 12 €/MWh, and $\psi^{-}_{t+k}$ is maintained at 4 €/MWh. This simplified setup is used to gain insight into the behavior of the proposed method, with a more practical case study deferred to Section 5. In this context, the optimal offering strategy is to use the 0.75 quantile of the wind power distributions. Specifically, this represents the independent optimal offering strategy for each WPP and the PM. Consistent with the recommendation of Bremnes, (2004), the optimal offers can be constructed using quantile regression. Moreover, since point forecasts—typically derived through mean regression—remain prevalent in real-world applications, we employ both mean and quantiles as base forecasts. A description of the cases is provided below.
@@
Base forecasts for both the aggregate and leaf levels are communicated as quantiles, derived from quantile regression models.
In this study, we evaluate various offering strategies, which include both independent and aggregate offering using reconciliation. The offering strategies under consideration are described as follows: (1) Independent offering: Each WPP submits offers based on their individual forecasts (i.e., the base forecasts). Thus, neither reconciliation nor the involvement of a PM is considered. We also include 3 aggregate offering strategies, in which the reconciled aggregate forecasts are used by the PM for offering. Concretely, the considered reconciliation approaches are described as follows. (2) Bottom-up reconciliation: The base forecasts at the leaf level are retained unchanged and then aggregated to generate the overall forecast. (3) Quality-oriented reconciliation: The base forecasts are reconciled within a learning-based model (17), where the loss function is defined as the mean squared error. (4) Value-oriented reconciliation: The base forecasts are reconciled using the learning-based model (17), with the training procedure outlined in Section 3. Particularly, the quality-based reconciliation is excluded in Case 2 because minimizing the total mean squared errors is irrelevant when the base forecasts are communicated as quantiles.
It is important to note that all of the previously discussed aggregate offering strategies employ the proportional allocation rule and set $\gamma_{i}$ as $\gamma_{i}^{\texttt{ge}}$, as outlined in Section 3.2. For both quality-oriented and value-oriented reconciliation models, we include the lagged wind power generation at both the leaf and aggregate levels as contextual features. To evaluate the performance of the offering strategies across a set of samples $\{(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t},\mathbf{y}_{t+k})\}_{t\in\mathcal{T}_{\texttt{test}}}$, we compute the average profit (AP) for each WPP over the test set. Specifically, for independent offering, using $\rho_{t+k}$ defined by (2), the average profit for WPP $i$ is given by
$\text{AP}_{i}=\frac{1}{|\mathcal{T}_{\texttt{test}}|}\sum_{t\in\mathcal{T}_{\texttt{test}}}\rho_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i}).$
In contrast, for portfolio-based aggregate offering, using $\rho_{t+k,i}^{(\texttt{AG})}$ from (16), the average profit for WPP $i$ is given by
$\text{AP}_{i}=\frac{1}{|\mathcal{T}_{\texttt{test}}|}\sum_{t\in\mathcal{T}_{\texttt{test}}}\rho_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k}).$

### v2

@@ 3.2.3 Profit Allocation within the Aggregation
Let $\tilde{y}_{t+k,\mathrm{sum}|t}$ be the offer by the PM for the trading hour $t+k$, and $\hat{y}_{t+k,1|t},\cdots,\hat{y}_{t+k,m|t}$} be a set of $m$ independent offers. Then, the extra profit is defined as
$R=\sum_{i=1}^{m}c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}}),$
where $c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})$ is the cost for the aggregate offer $\tilde{y}_{t+k,\mathrm{sum}|t}$ post-reconciliation, and $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$ is the cost for the independent offer using the base forecast $\hat{y}_{t+k,i|t}$.
@@
$\sum_{i=1}^{m}c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})=(1-w)\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})+wc_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}}).$
@@
$c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})\leq\sum_{i=1}^{m}c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}),$ (27)
where the inequality results from Proposition 3.1. Therefore, it can be seen that the designed cost allocation rule is not efficient, unless $w=1$, as we have
$\sum_{i=1}^{m}c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-\sum_{i=1}^{m}c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq R.$ (28)
In other words, the total extra profit allocated to WPPs is lower than the extra profit generated by aggregated trading. Specifically, the remaining extra profit is assigned to the PM as the payoff for their coordination efforts, denoted as $R^{\mathrm{pm}}$, i.e.,
$R^{\mathrm{pm}}=(1-w)\left[\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})-c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})\right].$ (29)
@@ 3.3 Model Training via Empirical Risk Minimization
Indeed, problem (24) can be viewed as a machine learning problem with constraints. As demonstrated by chamon2020probably, learning with constraints is, in theory, no more difficult than unconstrained learning, and empirical risk minimization remains an appropriate approach for such problems. Specifically, they proposed a primal–dual gradient algorithm for constrained learning. Building on this, we adapt their framework and derive the empirical Lagrangian function for problem (24), i.e.
$\begin{split}L(\bm{\theta},\bm{\mu})&=-\sum_{i}^{m}\log\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]\right]\\ &+\sum_{i}^{m}\mu_{i}\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+},\end{split}$ (30)
@@
$\nabla_{\mu_{i}}=\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+},$
where $\tilde{\mathbf{y}}_{t+k|t}=\mathbf{S}g(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t};\bm{\theta})$. Hence, the main challenge of problem (31) lies in the inner minimization. However, abundant evidence has shown that the gradient descent algorithm can learn good parameters for neural networks (shalev2014understanding). Subsequently, we can iteratively minimize the Lagrangian with respect to $\bm{\theta}$ while keeping $\bm{\mu}$ constant, followed by adjusting the dual variables using the obtained minimizer.
@@
  $\mu_{i}^{\eta}\leftarrow\mu_{i}^{\eta-1}+\nu_{i}\left[\frac{1}{|\mathcal{T}_{b}|}\sum_{t\in\mathcal{T}_{b}}\left[c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]\right]^{+}$
@@
We now analyze the convexity of the Lagrangian function. The convexity of the second component of $L(\bm{\theta},\bm{\mu})$ is straightforward, as $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is convex in $\tilde{\mathbf{y}}_{t+k|t}$. For the first component of $L(\bm{\theta},\bm{\mu})$, we note that $\log(\cdot)$ is convex and decreasing, while $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is concave in $\tilde{\mathbf{y}}_{t+k|t}$. Since the combination of a convex, non-increasing function with a concave function preserves convexity (boyd2004convex), it follows that the first term in $L(\bm{\theta},\bm{\mu})$ is also convex. Therefore, the Lagrangian function $L(\bm{\theta},\bm{\mu})$ is convex in $\tilde{\mathbf{y}}_{t+k|t}$, enabling the use of the stochastic gradient descent algorithm to update $\bm{\theta}$ (shalev2014understanding). The explicit form of $L^{(b)}(\bm{\theta},\bm{\mu})$ is deferred to the Appendix. Letting $\bm{\mu}^{\eta}$ and $\bm{\theta}^{\eta}$ denote the parameter estimates at the $\eta$ iteration, the iterative learning procedure is summarized in Algorithm 1.
Remark 3.1.
Note that the loss function is non-stationary. However, under Assumptions 3.3 and 3.4, Algorithm 1 is guaranteed to converge sublinearly, as established by Theorem 1 by besbes2015non.
@@ 4 Simulations
In this section, we present several illustrative examples to validate the proposed approach using data from the Wind Toolkit [^3] https://www.nrel.gov/grid/wind-toolkit.html. Specifically, we have collected hourly datasets from four nearby wind farms, with capacities of 1.7496 MW, 2.9646 MW, 3.3777 MW, and 2.5272 MW, covering the period from 2007 to 2008. In particular, we follow the framework of munoz2023online, which examines an hourly online forward market followed by a balancing market with a dual-price settlement mechanism. The lead time, $k$, for both the base forecast and forecast reconciliation process is set to 1. The dataset is split into two segments: the first 80% is used for model training, while the remaining 20% is reserved for out-of-sample forecasting validation. In the following, we will introduce the setup and present the simulation results.
@@ 4.1 Experimental Setups
In our simulation studies, we investigate a simplified scenario where $\pi_{t+k}^{\mathrm{F}}$ is fixed at 25 €/MWh, $\psi^{+}_{t+k}$ is set to 12 €/MWh, and $\psi^{-}_{t+k}$ is maintained at 4 €/MWh. This simplified setup is used to gain insight into the behavior of the proposed method, with a more practical case study deferred to Section 5. In this context, the optimal offering strategy is to use the 0.75 quantile of the wind power distributions. Specifically, this represents the independent optimal offering strategy for each WPP and the PM. Therefore, optimal offers can be constructed using quantiles, which represent a specific type of point forecast (gneiting2011quantiles). Moreover, as point forecasts derived from mean regression are still widely used in practice, we consider both means and quantiles as base forecasts in our framework. While WPPs may act strategically by explicitly communicating quantiles, we assume that the PM remains agnostic to the specific form of the base forecasts. A description of the cases is provided below.
@@
Base forecasts for both the aggregate and leaf levels are communicated as quantiles, derived from quantile regression models. [^4] Bottom-up and quality-oriented reconciliation methods are behaviorally plausible but decision-theoretically sub-optimal in this case, which is the tension we intend to highlight.
In this study, we evaluate various offering strategies, which include both independent and aggregate offering using reconciliation. The offering strategies under consideration are described as follows: (1) Independent offering: Each WPP submits offers based on their individual forecasts (i.e., the base forecasts). Thus, neither reconciliation nor the involvement of a PM is considered. We also include 3 aggregate offering strategies, in which the reconciled aggregate forecasts are used by the PM for offering. Concretely, the considered reconciliation approaches are described as follows. (2) Bottom-up reconciliation: The base forecasts at the leaf level are retained unchanged and then aggregated to generate the overall forecast. (3) Quality-oriented reconciliation: The base forecasts are reconciled within a learning-based model (19), where the loss function is defined as the mean squared error. (4) Value-oriented reconciliation: The base forecasts are reconciled using the learning-based model (19), with the training procedure outlined in Section 3. In addition, we establish linear counterparts for both quality-oriented and value-oriented reconciliation in Case 1, where the quality-oriented model reduces to the framework proposed by hyndman2011optimal.
It is important to note that all of the previously discussed aggregate offering strategies employ the proportional allocation rule and set $\gamma_{i}$ as $\gamma_{i}^{\mathrm{ge}}$, as outlined in Section 3.2. For both quality-oriented and value-oriented reconciliation models, we include the lagged wind power generation at both the leaf and aggregate levels as contextual features. To evaluate the performance of the offering strategies across a set of samples $\{(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t},\mathbf{y}_{t+k})\}_{t\in\mathcal{T}_{\mathrm{test}}}$, we compute the average profit (AP) for each WPP over the test set. Specifically, for independent offering, using $\rho_{t+k}$ defined by (2), the average profit for WPP $i$ is given by
$\mathrm{AP}_{i}=\frac{1}{|\mathcal{T}_{\mathrm{test}}|}\sum_{t\in\mathcal{T}_{\mathrm{test}}}\rho_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i}).$
In contrast, for portfolio-based aggregate offering, using $\rho_{t+k,i}^{(\mathrm{AG})}$ from (16), the average profit for WPP $i$ is given by
$\mathrm{AP}_{i}=\frac{1}{|\mathcal{T}_{\mathrm{test}}|}\sum_{t\in\mathcal{T}_{\mathrm{test}}}\rho_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k}).$